OpenStudy (anonymous):

Find the difference. (9/12)-(4/9)

OpenStudy (neer2890):

\[\frac{ 9 }{ 12 }-\frac{ 4 }{ 9 }=\frac{ 3 }{ 4 }-\frac{ 4 }{ 9 }=\frac{ 3*9-4*4 }{ 36 }=\frac{ 27-16 }{ 36 }=\frac{ 11 }{ 36 }\]